# Agentic Multi-Persona Framework for Evidence-Aware Fake News Detection

### v1

@@ Abstract.
The rapid proliferation of online misinformation poses significant risks to public trust, policy, and safety, necessitating reliable automated fake news detection. Existing methods often struggle with multimodal content, domain generalization, and explainability. We propose AMPEND-LS, an agentic multi-persona evidence-grounded framework with LLM–SLM synergy for multimodal fake news detection. AMPEND-LS integrates textual, visual, and contextual signals through a structured reasoning pipeline powered by LLMs, augmented with reverse image search, knowledge graph paths, and persuasion strategy analysis. To improve reliability, we introduce a credibility fusion mechanism combining semantic similarity, domain trustworthiness, and temporal context, and a complementary SLM classifier to mitigate LLM uncertainty and hallucinations. Extensive experiments across three benchmark datasets demonstrate that AMPEND-LS consistently outperformed state-of-the-art baselines in accuracy, F1 score, and robustness. Qualitative case studies further highlight its transparent reasoning and resilience against evolving misinformation. This work advances the development of adaptive, explainable, and evidence-aware systems for safeguarding online information integrity.
Fake News Detection, Misinformation Detection, Information Retrieval, LLM
$c$s: Computing methodologies Reasoning about belief and knowledgeccs: Information systems Information retrieval
1. Introduction
@@
News plays a central role in shaping public opinion, guiding decisions, and maintaining social stability. With digital media and social platforms accelerating global dissemination, information now spreads almost instantly. However, this connectivity also amplifies the reach of fake news (zhou2020survey), which can erode trust, cause confusion, and trigger severe consequences in health, politics, and public safety. These societal risks have made fake news detection a critical research area for safeguarding the integrity of online ecosystems (shu2020fakenewsnet; das2025fakenewsdetectionllm). Early work on fake news detection relied on manual fact-checking, reliable but unscalable (graves2018understanding). Automated approaches reframed detection as supervised classification, leveraging linguistic cues, user behavior, and propagation patterns (shu2017fake). Deep learning further advanced these models by capturing complex textual, semantic, and contextual dependencies (zhou2020survey). Yet, the rapid evolution of misinformation and heterogeneous news sources reveals persistent gaps, particularly in multimodal integration, cross-domain generalization, and reliance on resource-intensive annotations.
@@
Over time, methods have evolved from text-only classifiers to multimodal, domain-adaptive, and large language model (LLM)-driven paradigms. Multimodal approaches, from early joint encoders EANN (wang2018eann)/ MVAE (khattar2019mvae) to contrastive and graph-enhanced models (wei2025structure; 10.1145/3581783.3612423), demonstrate the value of heterogeneous cues, but suffer from fixed encoders, shallow fusion, and costly pipelines. Domain adaptation methods (WU2025111485; 10.1145/3664647.3681317) mitigate distribution shifts yet depend on domain metadata and remain brittle under unseen domains. More recently, LLM-based methods (hu2025synergizingllmsgloballabel; modzelewski2025pcotpersuasionaugmentedchainthought) enhance reasoning and semantic understanding, but face high computational costs, weak multimodal integration, and shallow, non-evidence-grounded explanations, limiting use in high-stakes settings. Collectively, existing approaches capture correlations but lack scalability, robustness, and evidence-grounding transparency (harris2024review).
@@
To address these challenges, we propose an explainable, evidence-grounded multi-persona agentic framework for multimodal fake news detection. Our framework integrates evidence retrieval through text-based web search and reverse image search, with each piece of evidence evaluated by a composite reliability score combining lexical and semantic similarity, domain credibility, and temporal consistency. In parallel, a knowledge graph (KG) encodes entity–relationship structures to provide factual grounding. These heterogeneous signals are synthesized by a multi-persona agentic module that engages in multi-round, persona-driven questioning to iteratively refine its contextual memory. Building on this, we design a modular classification pipeline where an LLM-based classifier produces pseudo labels with justifications, and a lightweight small language model (SLM) distills both the agent’s reasoning and LLM outputs. Importantly, for uncertain cases, we add persuasion-based analysis that detects manipulative strategies, computes a persuasion index, and refines reasoning. This design ensures low latency, high efficiency, and strong explainability, advancing multimodal fake news detection.
Our key contributions are summarized below:
@@
2. Related Work
@@
Fake news detection is a challenging classification task requiring reasoning over diverse information sources. Early methods struggled with multimodal evidence, while recent baselines integrate text and images but rely on simple fusion. Current advances can be broadly grouped into three categories: multimodal techniques (khattar2019mvae; wei2025structure), domain adaptation methods, and advanced language model–based approaches.
@@
(i) Multimodal Fake News Analysis: Early multimodal methods, such as EANN (wang2018eann) and MVAE (khattar2019mvae), combined textual and visual features but relied on fixed encoders with limited adaptability. Subsequent works tackled cross-modal inconsistencies via similarity- and ambiguity-aware fusion (SAFE (zhou2020similarity), CAFE (chen2022cross)), while SpotFake (singhal2019spotfake) integrated BERT for text yet retained shallow fusion. More recently, contrastive learning has dominated, with StruACL (wei2025structure), COOLANT (10.1145/3581783.3613850), and MFCL (CHEN2025112800) improving alignment through adversarial or contrastive objectives, and GAMED (10.1145/3701551.3703541) exploiting mixture-of-experts for feature discrimination. Graph-based approaches enhance this by incorporating semantic and knowledge graphs (HSEN (10.1145/3581783.3612423), KAMP (10.1145/3696410.3714532), GS2F (10.1145/3708536)). Transformer-based architectures further refine fusion with hierarchical attention (10.1145/3581783.3612423; 10.1145/3581783.3613850; 10.1145/3696410.3714532; li-etal-2025-imrrf; 10834537). Despite these advances, most methods still rely on pretrained encoders, incur high computational cost, and lack explicit evidence-grounded reasoning, explainability, and scalability, limiting robustness in high-stakes misinformation detection.
Figure 1. Proposed Framework: AMPEND-LS
@@
(ii) Domain Adaptation-based Methods: Domain adaptation is extensively studied to handle distribution shifts in fake news detection (wei-etal-2025-cross). Recent methods include hierarchical and transformer-based architectures such as HFGD (WU2025111485), MMDFND (10.1145/3664647.3681317), BREAK (10834537), and StruACL-TGN (wei2025structure), which leverage attention or structure-aware mechanisms for cross- and multi-domain settings. Transfer learning approaches like MMHT (10.1145/3696410.3714517) disentangle veracity-relevant from domain-specific features, while PFND (11027119) applies graph convolutional networks to combine content and fact-based evidence. Generative models such as MVAE (khattar2019mvae) and GSFND (tong2025generate) align multimodal features or augment training data. However, these methods often rely on domain labels or engagement signals, incur high cost, and lack interpretability, robustness, and evidence-grounded reasoning in mismatched or unseen domains.
@@
(iii) Advance Language Model-based techniques: LLMs have advanced fake news detection through reasoning-augmented and robustness-driven strategies. GLPN-LLM (hu2025synergizingllmsgloballabel) incorporates global label reasoning, while PCoT (modzelewski2025pcotpersuasionaugmentedchainthought) augments chain-of-thought prompting with persuasion cues. Robustness has been pursued via hallucination informed reasoning (IMRRF) (li-etal-2025-imrrf), adversarial style generation (AdStyle) (10.1145/3696410.3714569), and laundering-resistant detection (das2025fakenewsdetectionllm). Retrieval and collaboration approaches include self-reflective retrieval (SR3) (Zhou_Zhang_Tan_Zhang_Li_2025) and multi-agent LLM–SLM co-learning (10704605). Data augmentation methods such as GSFND (tong2025generate) and TripleFact (xu2025triplefact) employ generative sampling and evidence-grounded triples. Yet, LLM-based systems remain limited by high computational cost, shallow multimodal integration, vulnerability to data contamination, and difficulties adapting to rapidly evolving misinformation.
@@
Positioning of our Paper: To overcome the constraints of current approaches, we propose AMPEND-LS, an evidence-grounded, multi-persona agentic framework for robust fake news detection. Unlike domain adaptation methods that depend on domain labels or user engagement, and multimodal models limited by shallow fusion or costly encoders, AMPEND-LS explicitly integrates textual, visual, and contextual evidence with structured reasoning. A composite reliability score fuses lexical–semantic similarity, domain credibility, and temporal consistency, providing fine-grained veracity signals. Multi-persona agentic reasoning enables iterative claim validation through diverse expert roles, generating transparent justifications. This reasoning is distilled into a hybrid classification pipeline, where an LLM classifier captures rich contextual memory and a lightweight SLM provides efficient, uncertainty-aware verification. By combining evidence fusion, agentic reasoning, and scalable classification, AMPEND-LS achieves stronger factual grounding, interpretability, and adaptability than existing methods, positioning it as a deployable and trustworthy framework for combating misinformation in dynamic, multi-domain environments.
3. Problem Formulation
@@
4. Methodology
Our proposed framework, AMPEND-LS (Figure 1), integrates textual, visual, and contextual cues into a unified pipeline designed to (i) handle multimodal inputs, (ii) ground predictions in credible evidence, and (iii) provide explainable reasoning.
@@ 4.1. Preprocessing
4.2. Evidence and Facts Retrieval
@@
(1) $\scriptsize s_{1}=\sigma~\!\bigl(BM25(\mathcal{C},top\text{-}K_{1}(\mathcal{D}))\bigr),\qquad s_{2}=BS(\mathcal{C},top\text{-}K_{1}(\mathcal{D})),$
where, $s_{1}$ measures lexical relevance via BM25 (bm25), $s_{2}$ measures semantic similarity via BERT-Score (bert_score), and $\sigma$ is a normalization function. Formally, the evidence retrieval operator is defined as:
$f_{\text{retr}}:\mathcal{C}\;\mapsto\;\{(d_{i},s_{1}(d_{i}),s_{2}(d_{i}))\}_{i=1}^{K_{1}}.$
@@
(2) $\scriptsize s_{3}(d_{i})=\{\text{High}\!\mapsto\!\alpha_{h},\;\text{Medium}\!\mapsto\!\alpha_{m},\;\text{Unknown}\!\mapsto\!\alpha_{u},\;\text{Low}\!\mapsto\!\alpha_{l}\}.$
with ordering constraints: $\alpha_{h}>\alpha_{m}>\alpha_{u}>\alpha_{l},~~\alpha_{j}\in[0,1],\;j\in\{h,m,u,l\}.$ Typical settings are $\alpha_{h}\approx 1$, $\alpha_{l}\approx 0$, and $\alpha_{u}\approx 0.5$ to avoid unfair exclusion of less popular domains. The credibility operator is therefore defined as:
$f_{\text{cred}}:top\text{-}K_{1}(\mathcal{D})\;\mapsto\;\mathbf{s}_{3}=[s_{3}(d_{1}),\dots,s_{3}(d_{K_{1}})],$
which penalizes evidence from low-credibility sources while prioritizing reputable domains for claim verification.
@@
(3) $\scriptsize s_{4}(d_{i})=\begin{cases}\max(\alpha_{t},1-\Delta_{i}/365),&t_{i}\neq\text{unknown}\\ 0,&\text{otherwise}\end{cases}$
where, scores close to $1$ indicate strong temporal alignment with the current event. A lower bound of $\alpha_{t}$, a hyperparameter, prevents discarding slightly delayed but relevant reports, while documents without valid timestamps are excluded. The temporal filtering operator is thus defined as
$f_{\text{temp}}:top\text{-}{K_{1}}(\mathcal{D})\;\mapsto\;\mathbf{s}_{4}=[s_{4}(d_{1}),s_{4}(d_{2}),\dots,s_{4}(d_{K_{1}})].$
@@
(4) $\scriptsize\mathcal{R}(d_{i})=\lambda_{1}s_{1}(d_{i})+\lambda_{2}s_{2}(d_{i})+\lambda_{3}s_{3}(d_{i})+\lambda_{4}s_{4}(d_{i});\qquad\sum\nolimits_{j=1}^{4}\lambda_{j}=1$
@@
(v) Visual Evidence: Reverse image search is employed to trace the origin and context of images, enabling the retrieval of relevant external evidence to verify or refute potential fake news. We, therefore, perform reverse image search on the query image $\mathcal{I}$ using the Google Cloud Vision API. This yields a candidate set of documents with metadata: $\langle$ image URL, article URL, web entities, article title, article summary, date of publication $\rangle,$ from which irrelevant sources are discarded. For each candidate $d_{i}$, reliability is assessed using the same four scores as in textual evidence, lexical similarity, semantic similarity, domain credibility, and temporal consistency, resulting in an overall reliability score $\mathcal{R}(d_{i})$.
@@
4.3. Multi-step Multi-persona Agentic Framework (MMAF)
@@
Observation space: $\mathcal{O}=(\mathcal{N}_{e},\mathcal{K}_{\text{LLM}}),\quad\mathcal{N}_{e}=\{\mathcal{H},\mathcal{T}_{p},\mathcal{C},\mathcal{I}_{s},\mathcal{E}\},$ where, $\mathcal{H}$ = headline, $\mathcal{T}_{p}$ = preprocessed article text, $\mathcal{C}$ = extracted claim, $\mathcal{I}_{s}$ = image summary from image entities, $\mathcal{E}$ = retrieved evidence (credibility checks, reverse image search, KG facts), and $\mathcal{K}_{\text{LLM}}$ = LLM’s internal knowledge.
@@
(5) $\scriptsize\begin{split}q_{i}^{t}=\pi_{i}^{q}(o^{t},M^{t-1})\in\mathcal{Q}~;\hskip 16.38895pta_{i}^{t}=\pi^{a}(q_{i}^{t},o^{t},M^{t-1})\in\mathcal{A}_{\text{Ans}}~;\\ M^{t}=\mu(M^{t-1},(q_{i}^{t},a_{i}^{t})).\end{split}$
@@
4.4. LLM Pseudo-Labeling: Policy, Decision Rules, and Output

### v2

@@ Abstract
The rapid proliferation of online misinformation threatens the stability of digital social systems and poses significant risks to public trust, policy, and safety, necessitating reliable automated fake news detection. Existing methods often struggle with multimodal content, domain generalization, and explainability. We propose AMPEND-LS, an agentic multi-persona evidence-grounded framework with LLM–SLM synergy for multimodal fake news detection. AMPEND-LS integrates textual, visual, and contextual signals through a structured reasoning pipeline powered by LLMs, augmented with reverse image search, knowledge graph paths, and persuasion strategy analysis. To improve reliability, we introduce a credibility fusion mechanism combining semantic similarity, domain trustworthiness, and temporal context, and a complementary SLM classifier to mitigate LLM uncertainty and hallucinations. Extensive experiments across three benchmark datasets demonstrate that AMPEND-LS consistently outperformed state-of-the-art baselines in accuracy, F1 score, and robustness. Qualitative case studies further highlight its transparent reasoning and resilience against evolving misinformation. This work advances the development of adaptive, explainable, and evidence-aware systems for safeguarding online information integrity.
I Introduction
@@
Digital news ecosystems constitute a core component of modern computational social systems, influencing collective opinion formation, policy discourse, and societal stability. With the rapid expansion of online platforms, information now propagates across interconnected social networks at unprecedented speed and scale. While this connectivity enhances access to information, it simultaneously amplifies the spread of misinformation and fake news [43]. Such distortions of public discourse can undermine institutional trust, polarize communities, and generate harmful societal consequences in domains including public health, politics, and safety. Consequently, robust and transparent fake news detection has become a critical problem for safeguarding the integrity and resilience of digital social systems [25, 6]. Early mitigation efforts relied on manual fact-checking, which, although reliable, cannot scale to the velocity and volume of online information flows [7]. Automated approaches reframed detection as supervised classification, leveraging linguistic cues, user behavior, and propagation patterns [24]. Deep learning further advanced these models by capturing complex textual, semantic, and contextual dependencies [43]. Yet, the rapid evolution of misinformation and heterogeneous news sources reveals persistent gaps, particularly in multimodal integration, cross-domain generalization, and reliance on resource-intensive annotations.
@@
Over time, methods have evolved from text-only classifiers to multimodal, domain-adaptive, and large language model (LLM)-driven paradigms. Multimodal approaches, from early joint encoders EANN [30]/ MVAE [12] to contrastive and graph-enhanced models [31, 38], demonstrate the value of heterogeneous cues, but suffer from fixed encoders, shallow fusion, and costly pipelines. Domain adaptation methods [33, 27] mitigate distribution shifts yet depend on domain metadata and remain brittle under unseen domains. More recently, LLM-based methods [11, 17] enhance reasoning and semantic understanding, but face high computational costs, weak multimodal integration, and shallow, non-evidence-grounded explanations, limiting use in high-stakes settings. Collectively, existing approaches capture correlations but lack scalability, robustness, and evidence-grounding transparency [9].
@@
To address these challenges, we propose an explainable, evidence-grounded multi-persona agentic framework for multimodal fake news detection. Our framework integrates evidence retrieval through text-based web search and reverse image search, with each piece of evidence evaluated by a composite reliability score combining lexical and semantic similarity, domain credibility, and temporal consistency. In parallel, a knowledge graph (KG) encodes entity–relationship structures to provide factual grounding. These heterogeneous signals are synthesized by a multi-persona agentic module that engages in multi-round, persona-driven questioning to iteratively refine its contextual memory. Building on this, we design a modular classification pipeline where an LLM-based classifier produces pseudo labels with justifications, and a lightweight small language model (SLM) distills both the agent’s reasoning and LLM outputs. Importantly, for uncertain cases, we add persuasion-based analysis that detects manipulative strategies, computes a persuasion index, and refines reasoning. This design ensures low latency, high efficiency, and strong explainability, advancing multimodal fake news detection. Our key contributions are summarized below:
@@
The rest of this paper is organized as follows. Section II reviews related work. Section III introduces the problem formulation, and Section IV details the proposed methodology. Section V presents the experimental results and analysis. Limitations and future research directions are discussed in Section VI, followed by the conclusion in Section VII.
II Related Work
@@
Fake news detection is a challenging classification task requiring reasoning over diverse information sources. Early methods struggled with multimodal evidence, while recent baselines integrate text and images but rely on simple fusion. Current advances can be broadly grouped into three categories: multimodal techniques [12, 31], domain adaptation methods, and advanced language model–based approaches.
@@
(i) Multimodal Fake News Analysis: Early multimodal methods, such as EANN [30] and MVAE [12], combined textual and visual features but relied on fixed encoders with limited adaptability. Subsequent works tackled cross-modal inconsistencies via similarity- and ambiguity-aware fusion (SAFE [42], CAFE [5]), while SpotFake [26] integrated BERT for text yet retained shallow fusion. More recently, contrastive learning has dominated, with StruACL [31], COOLANT [29], and MFCL [2] improving alignment through adversarial or contrastive objectives, and GAMED [23] exploiting mixture-of-experts for feature discrimination. Graph-based approaches enhance this by incorporating semantic and knowledge graphs (HSEN [38], KAMP [40], GS2F [41]). Transformer-based architectures further refine fusion with hierarchical attention [38, 29, 40, 15, 37]. Despite these advances, most methods still rely on pretrained encoders, incur high computational cost, and lack explicit evidence-grounded reasoning, explainability, and scalability, limiting robustness in high-stakes misinformation detection.
Figure 1: Proposed Framework: AMPEND-LS
@@
(ii) Domain Adaptation-based Methods: Domain adaptation is extensively studied to handle distribution shifts in fake news detection [32]. Recent methods include hierarchical and transformer-based architectures such as HFGD [33], MMDFND [27], BREAK [37], and StruACL-TGN [31], which leverage attention or structure-aware mechanisms for cross- and multi-domain settings. Transfer learning approaches like MMHT [35] disentangle veracity-relevant from domain-specific features, while PFND [20] applies graph convolutional networks to combine content and fact-based evidence. Generative models such as MVAE [12] and GSFND [28] align multimodal features or augment training data. However, these methods often rely on domain labels or engagement signals, incur high cost, and lack interpretability, robustness, and evidence-grounded reasoning in mismatched or unseen domains.
@@
(iii) Advance Language Model-based techniques: LLMs have advanced fake news detection through reasoning-augmented and robustness-driven strategies. GLPN-LLM [11] incorporates global label reasoning, while PCoT [17] augments chain-of-thought prompting with persuasion cues. Robustness has been pursued via hallucination informed reasoning (IMRRF) [15], adversarial style generation (AdStyle) [18], and laundering-resistant detection [6]. Retrieval and collaboration approaches include self-reflective retrieval (SR3) [44] and multi-agent LLM–SLM co-learning [14]. Data augmentation methods such as GSFND [28] and TripleFact [34] employ generative sampling and evidence-grounded triples. Yet, LLM-based systems remain limited by high computational cost, shallow multimodal integration, vulnerability to data contamination, and difficulties adapting to rapidly evolving misinformation.
@@
Positioning of our Work: To overcome the constraints of current approaches, we propose AMPEND-LS, an evidence-grounded, multi-persona agentic framework for robust fake news detection. Unlike domain adaptation methods that depend on domain labels or user engagement, and multimodal models limited by shallow fusion or costly encoders, AMPEND-LS explicitly integrates textual, visual, and contextual evidence with structured reasoning. A composite reliability score fuses lexical–semantic similarity, domain credibility, and temporal consistency, providing fine-grained veracity signals. Multi-persona agentic reasoning enables iterative claim validation through diverse expert roles, generating transparent justifications. This reasoning is distilled into a hybrid classification pipeline, where an LLM classifier captures rich contextual memory and a lightweight SLM provides efficient, uncertainty-aware verification. By combining evidence fusion, agentic reasoning, and scalable classification, AMPEND-LS achieves stronger factual grounding, interpretability, and adaptability than existing methods, positioning it as a deployable and trustworthy framework for combating misinformation in dynamic, multi-domain environments.
III Problem Formulation
@@
IV Proposed Methodology
Our proposed framework, AMPEND-LS (Fig. 1), integrates textual, visual, and contextual cues into a unified pipeline designed to (i) handle multimodal inputs, (ii) ground predictions in credible evidence, and (iii) provide explainable reasoning.
@@ IV-A Preprocessing
IV-B Evidence and Facts Retrieval
@@
$\small s_{1}=\sigma~\bigl(BM25(\mathcal{C},top\text{-}K_{1}(\mathcal{D}))\bigr)~;~s_{2}=BS(\mathcal{C},top\text{-}K_{1}(\mathcal{D}));$ (1)
where, $s_{1}$ measures lexical relevance via BM25 [21], $s_{2}$ measures semantic similarity via BERT-Score [39], and $\sigma$ is a normalization function. Formally, the evidence retrieval operator is defined as: $f_{\text{retr}}:\mathcal{C}\mapsto\{(d_{i},s_{1}(d_{i}),s_{2}(d_{i}))\}_{i=1}^{K_{1}}.$
@@
$\scriptsize s_{3}(d_{i})=\{\text{High}\mapsto\alpha_{h},\text{Medium}\mapsto\alpha_{m},\text{Unknown}\mapsto\alpha_{u},\text{Low}\mapsto\alpha_{l}\}.$ (2)
with ordering constraints: $\alpha_{h}{>}\alpha_{m}{>}\alpha_{u}{>}\alpha_{l};~\alpha_{j}{\in}[0,1];~j\in\{h,m,u,l\}$. Typical settings are $\alpha_{h}{\approx}1$, $\alpha_{l}{\approx}0$, and $\alpha_{u}{\approx}0.5$ to avoid unfair exclusion of less popular domains. The credibility operator is therefore defined as: $f_{\text{cred}}:top\text{-}K_{1}(\mathcal{D})\mapsto\mathbf{s}_{3}=[s_{3}(d_{1}),\dots,s_{3}(d_{K_{1}})]$, which penalizes evidence from low-credibility sources while prioritizing reputable domains for claim verification.
@@
$\small s_{4}(d_{i})=\begin{cases}\max(\alpha_{t},1-\Delta_{i}/365);&t_{i}\neq\text{unknown}\\ 0;&\text{otherwise}\end{cases}$ (3)
where, scores close to $1$ indicate strong temporal alignment with the current event. A lower bound of $\alpha_{t}$, a hyperparameter, prevents discarding slightly delayed but relevant reports, while documents without valid timestamps are excluded. The temporal filtering operator is thus defined as $f_{\text{temp}}:top\text{-}{K_{1}}(\mathcal{D})\mapsto\mathbf{s}_{4}=[s_{4}(d_{1}),s_{4}(d_{2}),\dots,s_{4}(d_{K_{1}})].$
@@
$\scriptsize\mathcal{R}(d_{i})=\lambda_{1}s_{1}(d_{i})+\lambda_{2}s_{2}(d_{i})+\lambda_{3}s_{3}(d_{i})+\lambda_{4}s_{4}(d_{i});\quad\sum\nolimits_{j=1}^{4}\lambda_{j}=1$ (4)
@@
[…]
@@
IV-C Multi-step Multi-persona Agentic Framework (MMAF)
@@
Observation space: $\mathcal{O}=(\mathcal{N}_{e},\mathcal{K}_{\text{LLM}}),\quad\mathcal{N}_{e}=\{\mathcal{H},\mathcal{T}_{p},\mathcal{C},\mathcal{I}_{s},\mathcal{E}\}$, where $\mathcal{H}$ = headline, $\mathcal{T}_{p}$ = preprocessed article text, $\mathcal{C}$ = extracted claim, $\mathcal{I}_{s}$ = image summary from image entities, $\mathcal{E}$ = retrieved evidence (credibility checks, reverse image search, KG facts), and $\mathcal{K}_{\text{LLM}}$ = LLM’s internal knowledge.
@@
$\scriptsize\begin{split}q_{i}^{t}=\pi_{i}^{q}(o^{t},M^{t-1})\in\mathcal{Q}~;\hskip 16.38895pta_{i}^{t}=\pi^{a}(q_{i}^{t},o^{t},M^{t-1})\in\mathcal{A}_{\text{Ans}}~;\\ M^{t}=\mu(M^{t-1},(q_{i}^{t},a_{i}^{t})).\end{split}$ (5)
@@
IV-D LLM Pseudo-Labeling: Policy, Decision Rules, and Output